OpenStudy (nightmarenight):

Please help? I left my book at school... f(x) varies inversely with x and f(x) = –15 when x = 30. What is the inverse variation equation?

OpenStudy (anonymous):

\(y = \frac{k}{x}\) \(-15 = \frac{k}{30}\) \(k = -450\)

OpenStudy (anonymous):

\(y=\frac{k}{x}\) is the general form for your inverse variation so you just need to substitute \(k\) back.

OpenStudy (nightmarenight):

I don't know what numbers go where,,, @Richard_Feynman

OpenStudy (anonymous):

\(k\) is the constant of proportionality and \(x\) is your variable.